OpenStudy (anonymous):

Find the scalar equation of the plane that contains the parallel and distinct lines x=1, (y-3)/4=z/2 and x=5, (y+5)/2=(z-3)/1?

OpenStudy (anonymous):

Answer is 7x+2y-4z-13=0

OpenStudy (anonymous):

the direction vector of the 1º line is(0,4,4) or (0,1,1) to make it more easy the direction of 2º line is (0,2,1) To find the plane equation containing this 2 lines find the vector product of this 2 vectors, and later make that the plane pass through one of the points of the lines (1,3,0) or (5,-5,3)

OpenStudy (anonymous):

@Ruka

OpenStudy (anonymous):

Where did you get those direction vectors?

OpenStudy (anonymous):

(y-3)/4=z/4 this is line equation in continuous form

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

ok. P is point, V direction vector. Line equation: P(a,b,c) + t*V(i,j,k) = (x,y,z)

OpenStudy (anonymous):

this is equivalent to 3 equations: x=a+t*i y=b+t*j z=c+t*k solving all of them for t: x-a/i =t y-b/j=t z-c/k=t put them all together: x-a/i=y-b/j=z-c/k like you can see, in the denominators you have the components of the direction vector and in the nominator the points through which the lione goes

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Where did the other 4 come from? For the first one, shouldn't it be (0, 4, 2)? =(0, 2, 1) ?

OpenStudy (anonymous):

oh, sry i made a mistake, you right

OpenStudy (anonymous):

(0, 4, 2)=(0, 2, 1)

OpenStudy (anonymous):

Because the lines are parallel...

OpenStudy (anonymous):

you could take perpendicular vector to any of this two (0,-1,2) and find it's vector product with (0,2,1)

OpenStudy (anonymous):

or use the points through which this lines goes to find the other vector, which won't be paralel

OpenStudy (anonymous):

Huh? How do you get a, b, and c? ax+by+cz+d=0

OpenStudy (anonymous):

(5,-5,3) -(1,3,0) =(4,-8,3)

OpenStudy (anonymous):

Because the direction vector needs to equal (-7, -2, 4) or (7, 2, -4)

OpenStudy (anonymous):

ok. vector product of (0,2,1) x (4,-8,3)= (14,4,-8) plane equation using this point (5,-5,3): 14(x-5)+4(y+5)-8(z-3)=0 14x-70 +4y+20-8z+24=0 14x+4y-8z=70-20-24=26 which is you answer multiplied by 2

OpenStudy (anonymous):

(5,-5,3) -(1,3,0) =(4,-8,3)

OpenStudy (anonymous):

the 2 points of the lines

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):